OpenStudy (anonymous):

Solve 5x2 = –45x help!

OpenStudy (campbell_st):

let the equation equal zero by rewriting it \[5x^2 + 45x = 0\] remove a common factor of 5x \[5x(x + 9) = 0\] this should help you solve the quadratic

OpenStudy (anonymous):

5x^2 = -45x -> 5x^2 +45x = 0 So, x = 0, or x = -9.

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

No problem kiddo.